media%2F08b%2F08b14491-5c8e-41bc-8387-c2
An accelerometer can be modeled as a second order dv dv 2 dt Problem 3: system: dt the acceleration. m, c, and k are the physical properties o where v is the output voltage and a is m the accelerometer: mass, damping coefficient and stiffiness y. natural fiequency, damping ratio and sensitvity er (a) What of the

2018 A sinusoid has been sampled poorly. The result is shown belo could independently explain this misrepresentation e. w. List two errors that 0.3 0.7 1.2 1.6 Time (sec)
Answer #1

m\frac{d^2 v}{dt^2}+c\frac{dv}{dt}+k v = Ba

Laplace transform:

(ms^2+cs+k) V(s) = B A(s)

\frac {V(s)} {A(s)}= \frac{B} {(ms^2+cs+k)}

\frac {V(s)} {A(s)}= \frac{B} {m}*\frac{1}{(s^2+(c/m)s+(k/m))}

roots of the quadratic equation in the denominator are:

\frac{-(c/m)\pm \sqrt ((c/m)^2 - 4(k/m))}{2}

-(c/2m)\pm \sqrt(c^2/4m^2 - (k/m)

w_n = \sqrt(k/m) - Natural frequency

2\zeta w_n = c/m\Rightarrow \zeta = \frac{c}{2\sqrt(km)} - Damping ratio

b and c )

07wn J 26